OpenStudy (anonymous):

The probability that Shean and Nida pass an examination are 3/4 and 3/5 respectively. Find the probability of both boys failing the examination.

OpenStudy (anonymous):

let probability of Shean pass is p(s)=3/4 similarly Nida pass is p(n)=3/5 then for fail means nothing but negation/inverse of pass that is 1-p(s) and 1-p(n)

OpenStudy (anonymous):

The probability that shean fails is 1/4 and the probability that Nida fails is 2/5 because they only tell you the probability that they pass. now \[P(AandB)\] is\[P(A)*P(B)\]

OpenStudy (anonymous):

@chetan552, @mihirb. Logical solutions, but got confused to when I saw the listed options.

OpenStudy (anonymous):

at last have you clarified or not...?? if not tell me the options..

OpenStudy (anonymous):

1/3,3/10,9/20,1/3

OpenStudy (anonymous):

then i am sorry... i think answer must be 1/10...

OpenStudy (anonymous):

Thank you guys. Wanted to be sure I wasn't wrong.